OpenStudy (hippoharry):

What number must you add to complete the square? x2 + 4x = 3 A. 4 B. 8 C. 2 D. 16

OpenStudy (anonymous):

hmmm do you know what a "perfect square trinomial" is?

OpenStudy (anonymous):

or have you covered that yet?

OpenStudy (hippoharry):

Definition of a Perfect Square Binomial. A perfect square binomial is a trinomial that when factored gives you the square of a binomial. @jdoe0001

OpenStudy (anonymous):

hmmm have you covered it yet? or is that from google?

OpenStudy (hippoharry):

i have covered it. but its just confusing what the question they are asking @jdoe

OpenStudy (anonymous):

well let's take a peek at \(\bf x^2 + 4x = 3\) and do some grouping first so, we end up with \(\bf x^2 + 4x = 3\implies (x^2+4x+{\color{red}{ \square ?}}^2)=3\) so.. what's our missing value there, to get a perfect square inside the parentheses anyway?